Peoples Limited has the following securities in issue:



• 1 million ordinary shares currently trading at N$1.30 each. The last dividend paid



amounted to 15.5 cents and the earnings and dividends of Peoples are expected



to grow at 4% per annum.



• 250 000 non-redeemable 7% preference shares with a nominal value of



N$1.10 each. These shares are currently trading at N$0.65 per share.



• 1 550 12% unsecured bonds with a nominal value of N$100 each



currently trading at N$85 each. These bonds are redeemable in 3 years’



time at the nominal value.



Required:



(a) Calculate the weighted average cost of capital for Peoples Limited using market



values. (10 marks)



(b) Is there another way to calculate the weighted average cost of capital?



Support your answer with a brief explanation. Define all terms used in



your answer.

1
Expert's answer
2022-03-21T00:23:55-0400

(a)capital:

find the total capital. To do this, we add our own and borrowed capital


"1 000 000\times1.3+250 000\times1.3+1550\times100=1 730 000"


Find the rates of return on equity through the Gordon model

"r1=\frac{0.155(1+0.04)}{1.3}+0.04=0.1640"


"d2=\frac{0.65}{1.1}=0.59"


"r2=\frac{0.59(1+0.04)}{1.1}+0.04=0.5978"


Find the rate of return on a bond

"i=\frac{\frac{100-85}{3}+12}{\frac{100+85}{2}}=18.38"

we substitute the obtained data into the original formula:


"WACC=\frac{1 300 000}{1 730 000}\times16.40+\frac{275000}{1 730 000}\times59.78+\frac{155 000}{1 730 000}\times18.38=23.47"


(b)It is believed that this is the only formula for calculating the weighted average formula.

But also calculate through the sum of all market values of all outstanding securities multiplied by the required rate of return for collateral divided by the sum of all market values of all outstanding securities.

Also, the WACC formula can include tax effects, i.e. multiply the amount owed by one minus the tax rate.